# Truth in the Few: High-Value Data Selection for Efficient Multi-Modal Reasoning

### v1

@@ 3.2 Attention Confidence Estimator
3.3 Difficulty-aware Replacement Module
@@
where $k$ is the number of easy samples. This formulation identifies the top-$k$ most challenging samples according to the difficulty metric $D_{iff}^{i}$, while excluding those that do not meet the specified criteria. Note that the hardest samples, i.e., $D_{i}=1$, would be neglected, as they are demonstrated to be meaningless for training hintgrpo . This DRM can enhance the upper bound of the model’s ability to handle complex tasks, without introducing data redundancy and training costs.
@@ 4 Experiments
Training dataset. Main results in the Table 1 are based on models trained with the MM-Eureka dataset MM-Eureka , a high-quality multi-modal dataset for mathematical reasoning. To further validate the generalization of our RAP method, we evaluate models trained on the subset of Mulberry-260k dataset mulberry , a multi-modal learning-to-reason-and-reflect dataset.
@@
Evaluation. Similar to mulberry ; MM-Eureka , we evaluate models on both mathematical and general multi-modal reasoning tasks using the $pass@1$ metric, where $pass@1$ measures the percentage of problems correctly solved on the first attempt, under a zero-shot setting. For mathematical reasoning, we assess the model’s ability on four benchmarks: MathVista mathvista , MMStar mmstar , MathVerse mathverse , and WeMath wemath . For universal multi-modal reasoning, we evaluate on MMVet mmvet and LogicVista logicvista .
@@
Implementation details. Following prior methods MM-Eureka ; mulberry , we conduct our primary experiments on mathematical reasoning tasks, employing Qwen2.5-VL-3B and Qwen2.5-VL-7B qwen2025qwen25technicalreport as baseline models. First, we apply RAP to select cognitive samples using the initial model without any training. These samples are then used to train models within the EasyR1 easyr1 , employing the AdamW optimizer with a learning rate of 1e-6. Full-data training requires 1 epoch (107 steps), all others use 40 training steps with a batch size of 512 across 8 GPUs. The system prompt is available at our supplementary material. For accelerated generation in GRPO guo2025deepseek , we utilize the vLLM package kwon2023efficient . Finally, we set the hyperparameters $\sigma$ to 2.0, $\lambda_{c}$ and $\lambda_{a}$ to 0.5 and 0.1 for the CDE and ACE, respectively.
Table 1: Comparison with state-of-the-art methods. Experiments are conducted using the Qwen2.5-VL-3b qwen2025qwen25technicalreport and Qwen2.5-VL-7b qwen2025qwen25technicalreport , employing GRPO as the RL method. “Time” denotes the total computation cost, including data selection and training. Bold font denotes the best result.
Qwen2.5-VL-7b - - 68.70 56.07 39.31 35.90 59.13 44.52 50.61
Qwen2.5-VL-7b-s1 s1 (2025) 1,000 55.9 68.50 61.80 45.79 35.05 61.05 45.86 53.01
Qwen2.5-VL-7b-LIMO limo (2025) 4,093 111.9 69.90 61.33 45.74 34.67 59.08 44.74 52.58
Qwen2.5-VL-7b-LIMR limr (2025) 8,136 122.0 71.10 62.12 48.02 41.21 62.86 45.81 55.19
Qwen2.5-VL-3b - - 61.30 54.46 9.01 21.62 51.81 39.59 39.63
Qwen2.5-VL-3b-s1 s1 (2025) 1,000 38.4 62.60 54.53 37.41 27.52 53.02 39.59 45.78
Qwen2.5-VL-3b-LIMO limo (2025) 2,679 120.5 61.80 54.73 35.33 24.48 53.12 39.14 44.77
Qwen2.5-VL-3b-LIMR limr (2025) 21,303 60.9 63.10 54.66 35.43 26.76 53.25 40.95 45.69
4.1 Overall Comparison Results
@@
Comparing methods. We compare our approach with existing data selection methods, including: 1) s1 s1 , which utilizes the large-scale MLLMs to identify high-quality data; 2) LIMO limo that designs a difficulty-aware selection method to identify crucial samples; and 3) LIMR limr , which employs learning impact measurement to select a subset of training samples. Moreover, we also evaluate models trained on the full dataset (Full) as the baseline.
@@
Comparisons with state-of-the-art methods. We compare our RAP model with the latest data-efficient methods on six diverse datasets, including LIMO limo , s1 s1 and LIMR limr . The results shown in Table 1 reveal several key findings: 1) RAP consistently outperforms models trained with full corpora on all datasets. Remarkably, these improvements are achieved using only 9.5% or 7.9% of training data while reducing training time by 43% or 31%, supporting our hypothesis “truth in the few” that selected cognitive samples can achieve more effective multi-modal reasoning. 2) Moreover, the RAP method demonstrates a significant improvement of 7.33% and 6.95% over the LIMO limo and s1 s1 methods on WeMath wemath , respectively, which rely on data quality and manual selection. Such results indicate that focusing on the potential of each sample to improve multi-modal reasoning.
@@
Table 2: Comparison with state-of-the-art methods using Qwen2.5-VL-7b qwen2025qwen25technicalreport with the RL method RLOO rloo , evaluating our RAP method on different training datasets and RL algorithms.
@@
Qwen2.5-VL-7b-s1 s1 68.50 59.96 35.06 54.51
Qwen2.5-VL-7b-LIMO limo 68.80 60.11 35.23 54.71
Qwen2.5-VL-7b-LIMR limr 68.90 60.71 36.74 55.45
@@
Table 3: Comparison with state-of-the-art methods using InternVL3-2b internvl with the RL method GRPO guo2025deepseek , evaluating our RAP method across different model architectures.
@@
InternVL3-2b-s1 s1 56.80 60.47 11.63 42.97
InternVL3-2b-LIMO limo 56.70 59.82 11.92 42.81
InternVL3-2b-LIMR limr 57.10 61.33 12.44 43.62
@@
Effectiveness of RAP on different base models. As shown in Table 3, our method consistently surpasses other recent data selection methods when applied to the base model InternVL3-2b internvl . This outcome highlights the broad applicability and generalizability of the RAP framework, as the introduced CDE and ACE components effectively select training samples that activate the model’s multi-modal reasoning capability. Crucially, these components do not rely on exploiting model-specific inductive biases, thereby ensuring RAP’s adaptability on a wide range of model architectures. More results are available in our supplementary material.
Effectiveness of RAP for various RL methods and training datasets. To further validate the generalizability of RAP, Table 3 presents results using the Qwen2.5-VL-7B model, trained under two different configurations: 1) the RLOO RL paradigm rloo , and 2) the reduced Mulberry-10K dataset, a subset of Mulberry-260K mulberry . Despite these variations, RAP maintains consistent superiority compared to other methods, suggesting its generalization to different RL strategies and training datasets. We attribute this robustness to cognitive samples that facilitate genuine multi-modal reasoning rather than simply fitting data distributions.
4.2 Further Analysis
Ablation study. As presented in Table 4, we list the following conclusions: 1) The comparison between No.0 and No.3 indicates that integrating CDE and ACE can improve multi-modal reasoning in MLLMs. These results underscore the efficacy of RAP in eliminating language-prior biased and attention-biased samples, thereby enabling models to focus on essential “cognitive samples” and improving reasoning performance. 2) Comparing No.1 with No.3 shows that only using the CDE can improve performance, but worse than the full RAP. We hypothesize the reason is that while CDE identifies critical samples by detecting output discrepancies, it overlooks the intermediate reasoning process.
Table 4: Ablation study of RAP using Qwen2.5-VL-7B, trained using GRPO method for 40 steps.
@@ 4.2 Further Analysis
3) Moreover, the comparison between No.2 and No.3 demonstrates that the RCC can further refine the reasoning performance by replacing easy samples with more appropriate hard samples. Such improvements are due to that the DRM addresses a limitation of the CDE and ACE approaches, i.e., the tendency to retain simpler instances while neglecting challenging yet informative samples.
Comparison on cross-modal reasoning utilization. We evaluate the effectiveness of cognitive samples in enhancing cross-modal reasoning utilization on multi-modal reasoning tasks. We define cross-modal reasoning utilization as the proportion of instances in the MMStar dataset where the model correctly uses multi-modal inputs to answer questions, but fails when relying solely on textual inputs. As shown in Figure 5(c), models trained with cognitive samples show superior integration of cross-modal information, compared to baseline and the latest method LIMR limr . Such results highlight that the proposed CDE can effectively reduce models’ reliance on superficial linguistic prior by filtering out training samples exhibiting excessive language bias.
@@
Hence, models are encouraged to discover the relationship between image and text, thus improving the cross-modal reasoning utilization.
Comparison on cross-model generalization. To examine the cross-model generalizability of cognitive samples identified by RAP, we evaluate whether cognitive samples obtained using the Qwen2.5-VL-3b are useful for improving the reasoning of a distinctly structured model, InternVL3-2b, and vice versa. Comparative results presented in Figure 4 demonstrate that cognitive samples selected by our RAP method outperform the latest method LIMR limr , confirming the generalization of RAP in enhancing multi-modal reasoning for varying model frameworks. Note that we exclude comparisons with LIMO limo due to its similarity with s1 s1 .
Analysis on hyperparameter sensitivity. We further investigate the sensitivity of the hyperparameters $\lambda_{c}$ and $\lambda_{a}$ employed within the CDE and ACE components, respectively. As shown in Figure 5(a), we visualize the distribution of output discrepancies between multi-modal and text-only inputs on the full MM-Eureka dataset, revealing a significant presence of samples where multi-modal reasoning is not necessary to solve the task. Experimental results depicted in Figure 5(b) suggest that optimal performance is achieved with $\lambda_{c}=0.1$ and $\lambda_{a}=0.5$. Performance degradation is observed when both parameters fall below these optimal thresholds, with values lower than 0.1 for $\lambda_{c}$ and 0.5 for $\lambda_{a}$ leading to significant deterioration. This decline is attributed to the inherently uneven distribution of the output-level discrepancy and attention confidence, which are concentrated at the extremes.
Qualitative analysis. We provide a qualitative analysis by visualizing cognitive samples and comparing the reasoning processes of our RAP with the latest LIMR method as follows: 1) Visualization of cognitive samples. As shown in Figure 6(a), we present a typical example of the cognitive samples selected by our RAP method, where the model effectively integrates both geometric information and logical reasoning to infer the area. This case suggests that the cognitive samples selected by our RAP method exhibit two important characteristics: (a) the necessary of multi-modal information, as evidenced by the significant discrepancy between reasoning outcomes using multi-modal and text-only inputs; and (b) the avoidance of overemphasis on irrelevant or meaningless tokens, ensuring that the model focuses on the most informative features for accurate reasoning. 2) Comparison case analysis. As shown in Figure 6(b), we compare the reasoning process of our RAP method with that of the state-of-the-art LIMR approach limr . For example, LIMR fails to integrate cross-modal information, leading to incorrect computations of both the central angle and the inscribed angle. In contrast, the model trained using samples selected by RAP correctly applies geometric principles and multi-modal integration to arrive at the correct solution. These comparisons underscore the advantage of training with cognitive samples derived through our RAP method, which enables the model to effectively leverage multi-modal information, resulting in more reliable reasoning outcomes.
@@
Figure 6: Illustration of (a) characteristics of cognitive samples selected by our RAP method and (b) Comparison of the reasoning processes between our RAP method and the state-of-the-art LIMR limr .
4.3 Key Insights and Discussion
Effect of RAP on reasoning upper bound. As illustrated in Figure 7(a), our RAP method converges faster than the baseline, achieving optimal performance within 40 training steps compared to 100 for the baseline. These results demonstrate RAP’s efficiency in enhancing multi-modal reasoning while reducing training overhead and data redundancy. However, further analysis reveals differences between the full RAP model and its variant using only ACE and CDE, particularly during later training stages. This occurs because ACE and CDE inevitably retain easy examples, while discarding challenging yet informative samples. For example, multi-modal examples predicted correctly in only one out of five attempts but consistently mispredicted under text-only conditions may be mistakenly eliminated when the discrepancy threshold $\lambda_{c}$ exceeds 0.2. Such filtering reduces dataset complexity, limiting the model’s overall reasoning upper bound. To address this, we propose the Difficulty-aware Replacement Module (DRM) to explicitly replace easy samples with these informative yet challenging examples, thus elevating reasoning performance in later training stages.

### v2

@@ III-B Attention Confidence Estimator
TABLE I: Comparison with state-of-the-art methods. Experiments are conducted using the Qwen2.5-VL-3b [1] and Qwen2.5-VL-7b [1], employing GRPO as the RL method. “Time” denotes the total computation cost, including data selection and training. Bold font denotes the best result.
Qwen2.5-VL-7b - - 68.70 56.07 39.31 35.90 59.13 44.52 50.61
Qwen2.5-VL-7b-s1 [8] (2025) 1,000 55.9 68.50 61.80 45.79 35.05 61.05 45.86 53.01
Qwen2.5-VL-7b-LIMO [12] (2025) 4,093 111.9 69.90 61.33 45.74 34.67 59.08 44.74 52.58
Qwen2.5-VL-7b-LIMR [9] (2025) 8,136 122.0 71.10 62.12 48.02 41.21 62.86 45.81 55.19
Qwen2.5-VL-3b - - 61.30 54.46 9.01 21.62 51.81 39.59 39.63
Qwen2.5-VL-3b-s1 [8] (2025) 1,000 38.4 62.60 54.53 37.41 27.52 53.02 39.59 45.78
Qwen2.5-VL-3b-LIMO [12] (2025) 2,679 120.5 61.80 54.73 35.33 24.48 53.12 39.14 44.77
Qwen2.5-VL-3b-LIMR [9] (2025) 21,303 60.9 63.10 54.66 35.43 26.76 53.25 40.95 45.69
III-C Difficulty-aware Replacement Module
@@
where $k$ is the number of easy samples. This formulation identifies the top-$k$ most challenging samples according to the difficulty metric $D_{iff}^{i}$, while excluding those that do not meet the specified criteria. Note that the hardest samples, i.e., $D_{i}=1$, would be neglected, as they are demonstrated to be meaningless for training [35]. This DRM can enhance the upper bound of the model’s ability to handle complex tasks, without introducing data redundancy and training costs.
@@ IV Experiments
Training dataset. Main results in the Table I are based on models trained with the MM-Eureka dataset [26], a high-quality multi-modal dataset for mathematical reasoning. To further validate the generalization of our RAP method, we evaluate models trained on the subset of Mulberry-260k dataset [25], a multi-modal learning-to-reason-and-reflect dataset.
@@
Evaluation. Similar to [25, 26], we evaluate models on both mathematical and general multi-modal reasoning tasks using the $pass@1$ metric, where $pass@1$ measures the percentage of problems correctly solved on the first attempt, under a zero-shot setting. For mathematical reasoning, we assess the model’s ability on four benchmark datasets: MathVista [44], MMStar [45], MathVerse [46], and WeMath [47]. For universal multi-modal reasoning, we evaluate on MMVet [48] and LogicVista [49].
@@
Implementation details. Following prior methods [26, 25], we conduct our primary experiments on mathematical reasoning tasks, employing Qwen2.5-VL-3B and Qwen2.5-VL-7B [1] as baseline models. First, we apply RAP to select cognitive samples using the initial model without any training. These samples are then used to train models within the EasyR1 [50], employing the AdamW optimizer with a learning rate of 1e-6. Full-data training requires 1 epoch (107 steps), all others use 40 training steps with a batch size of 512 across 8 GPUs. The system prompt is available at our supplementary material. For accelerated generation in GRPO [3], we utilize the vLLM package [51]. Finally, we set the hyperparameters $\sigma$ to 2.0, $\lambda_{c}$ and $\lambda_{a}$ to 0.5 and 0.1 for the CDE and ACE, respectively.
IV-A Overall Comparison Results
@@
Comparing methods. We compare our approach with existing data selection methods, including: 1) s1 [8], which utilizes the large-scale MLLMs to identify high-quality data; 2) LIMO [12] that designs a difficulty-aware selection method to identify crucial samples; and 3) LIMR [9], which employs learning impact measurement to select a subset of training samples. Moreover, we also evaluate models trained on the full dataset (Full) as the baseline.
@@
Comparisons with state-of-the-art methods. We compare our RAP model with the latest data-efficient methods on six diverse datasets, including LIMO [12], s1 [8] and LIMR [9]. The results shown in Table I reveal several key findings: 1) RAP consistently outperforms models trained with full corpora on all datasets. Remarkably, these improvements are achieved using only 9.5% or 7.9% of training data while reducing training time by 43% or 31%, supporting our hypothesis “truth in the few” that selected cognitive samples can achieve more effective multi-modal reasoning. 2) Moreover, the RAP method demonstrates a significant improvement of 7.33% and 6.95% over the LIMO [12] and s1 [8] methods on WeMath [47], respectively, which rely on data quality and manual selection. Such results indicate that focusing on the potential of each sample to improve multi-modal reasoning.
@@
TABLE II: Comparison with state-of-the-art methods using Qwen2.5-VL-7b [1] with the RL method RLOO [21], evaluating our RAP method on different training datasets and RL algorithms.
@@
Qwen2.5-VL-7b-s1 [8] 68.50 59.96 35.06 54.51
Qwen2.5-VL-7b-LIMO [12] 68.80 60.11 35.23 54.71
Qwen2.5-VL-7b-LIMR [9] 68.90 60.71 36.74 55.45
@@
TABLE III: Comparison with state-of-the-art methods using InternVL3-2b [52] with the RL method GRPO [3], evaluating our RAP method across different model architectures.
@@
InternVL3-2b-s1 [8] 56.80 60.47 11.63 42.97
InternVL3-2b-LIMO [12] 56.70 59.82 11.92 42.81
InternVL3-2b-LIMR [9] 57.10 61.33 12.44 43.62
@@
Effectiveness of RAP on different base models. As shown in Table II, our method consistently surpasses other recent data selection methods when applied to the base model InternVL3-2b [52]. This outcome highlights the broad applicability and generalizability of the RAP framework, as the introduced CDE and ACE components effectively select training samples that activate the model’s multi-modal reasoning capability. Crucially, these components do not rely on exploiting model-specific inductive biases, thereby ensuring RAP’s adaptability on a wide range of model architectures. More results are available in our supplementary material.
Effectiveness of RAP for various RL methods and training datasets. To further validate the generalizability of RAP, Table III presents results using the Qwen2.5-VL-7B model, trained under two different configurations: 1) the RLOO RL paradigm [21], and 2) the reduced Mulberry-10K dataset, a subset of Mulberry-260K [25]. Despite these variations, RAP maintains consistent superiority compared to other methods, suggesting its generalization to different RL strategies and training datasets. We attribute this robustness to cognitive samples that facilitate genuine multi-modal reasoning rather than simply fitting data distributions.
TABLE IV: Ablation study of RAP using Qwen2.5-VL-7B, trained using GRPO method for 40 steps.
@@
IV-B Further Analysis
Ablation study. As presented in Table IV, we list the following conclusions: 1) The comparison between No.0 and No.3 indicates that integrating CDE and ACE can improve multi-modal reasoning in MLLMs. These results underscore the efficacy of RAP in eliminating language-prior biased and attention-biased samples, thereby enabling models to focus on essential “cognitive samples” and improving reasoning performance. 2) Comparing No.1 with No.3 shows that only using the CDE can improve performance, but worse than the full RAP. We hypothesize the reason is that while CDE identifies critical samples by detecting output discrepancies, it overlooks the intermediate reasoning process. 3) Moreover, the comparison between No.2 and No.3 demonstrates that the RCC can further refine the reasoning performance by replacing easy samples with more appropriate hard samples. Such improvements are due to that the DRM addresses a limitation of the CDE and ACE approaches, i.e., the tendency to retain simpler instances while neglecting challenging yet informative samples.
@@ IV-B Further Analysis
Comparison on cross-modal reasoning utilization. We evaluate the effectiveness of cognitive samples in enhancing cross-modal reasoning utilization on multi-modal reasoning tasks. We define cross-modal reasoning utilization as the proportion of instances in the MMStar dataset where the model correctly uses multi-modal inputs to answer questions, but fails when relying solely on textual inputs. As shown in Fig. 5(c), models trained with cognitive samples show superior integration of cross-modal information, compared to baseline and the latest method LIMR [9]. Such results highlight that the proposed CDE can effectively reduce models’ reliance on superficial linguistic prior by filtering out training samples exhibiting excessive language bias. Hence, models are encouraged to discover the relationship between image and text, thus improving the cross-modal reasoning utilization.
Comparison on cross-model generalization. To examine the cross-model generalizability of cognitive samples identified by RAP, we evaluate whether cognitive samples obtained using the Qwen2.5-VL-3b are useful for improving the reasoning of a distinctly structured model, InternVL3-2b, and vice versa. Comparative results presented in Fig. 4 demonstrate that cognitive samples selected by our RAP method outperform the latest method LIMR [9], confirming the generalization of RAP in enhancing multi-modal reasoning for varying model frameworks. Note that we exclude comparisons with LIMO [12] due to its similarity with s1 [8].
Analysis on hyperparameter sensitivity. We further investigate the sensitivity of the hyperparameters $\lambda_{c}$ and $\lambda_{a}$ employed within the CDE and ACE components, respectively. As shown in Fig. 5(a), we visualize the distribution of output discrepancies between multi-modal and text-only inputs on the full MM-Eureka dataset, revealing a significant presence of samples where multi-modal reasoning is not necessary to solve the task. Experimental results depicted in Fig. 5(b) suggest that optimal performance is achieved with $\lambda_{c}=0.1$ and $\lambda_{a}=0.5$. Performance degradation is observed when both parameters fall below these optimal thresholds, with values lower than 0.1 for $\lambda_{c}$ and 0.5 for $\lambda_{a}$ leading to significant deterioration. This decline is attributed to the inherently uneven distribution of the output-level discrepancy and attention confidence, which are concentrated at the extremes.
Analysis on the choice of attention layers. To clarify our rationale for using last-layer attention, we provide additional experiments exploring attention maps from other layers. As illustrated in Section III-B and Fig. 2(b), attention-biased distributions typically manifest as excessively persistent high attention allocation across subsequent token sequences. Existing research [28] has demonstrated that attention distributions in the final layer effectively capture causal reasoning behaviors. Such findings intuitively highlight the critical role of the last-layer attention maps in evaluating the reliability of the model’s reasoning behavior. Inspired by this, we initially chose the last-layer attention distributions to effectively capture globally biased attention patterns that compromise genuine multi-modal reasoning.
However, we have conducted preliminary experiments to investigate the potential of attention confidence computed from intermediate layers. Recent work [53] has indicated layers 5 to 8 within large language models play a critical role in multi-modal fusion. Therefore, we selected the 6th transformer layer’s attention distribution as an alternative input for our ACE module and compared its efficacy against our original last-layer choice. The results presented in Table V demonstrate slight performance degradation when utilizing intermediate (6th-layer) attention maps. For most datasets, employing attention maps from the last layer consistently yielded superior performance. These results justify our choice of using last-layer attention maps.
TABLE V: Performance comparison between RAP and RAP-mid using QwenVL2.5-3b with the GRPO RL algorithm across three benchmarks. RAP-mid refers to the variant that utilizes intermediate (6th-layer) attention maps.
Method MathVista MMVet LogicVista WeMath Avg.
QwenVL2.5-3b-RAP-mid 64.60 54.29 41.74 29.17 47.45
Qwen2.5VL-3b-RAP (Ours) 64.90 54.63 41.61 29.33 47.62
Qualitative analysis. We provide a qualitative analysis by visualizing cognitive samples and comparing the reasoning processes of our RAP with the latest LIMR method [9] as follows: 1) Visualization of cognitive samples. As shown in Fig. 6(a), we present a typical example of the cognitive samples selected by our RAP method, where the model effectively integrates both geometric information and logical reasoning to infer the area. This case suggests that the cognitive samples selected by our RAP method exhibit two important characteristics: (a) the necessary of multi-modal information, as evidenced by the significant discrepancy between reasoning outcomes using multi-modal and text-only inputs; and (b) the avoidance of overemphasis on irrelevant or meaningless tokens, ensuring that the model focuses on the most informative features for accurate reasoning. 2) Comparison case analysis. As shown in Fig. 6(b), we compare the reasoning process of our RAP method with that of the state-of-the-art LIMR approach [9]. For example, LIMR fails to integrate cross-modal information, leading to incorrect computations of both the central angle and the inscribed angle. In contrast, the model trained using samples selected by RAP correctly applies geometric principles and multi-modal integration to arrive at the correct solution. These comparisons underscore the advantage of training with cognitive samples derived through our RAP method, which enables the model to effectively leverage multi-modal information, resulting in more reliable reasoning outcomes.
@@
Figure 6: Illustration of (a) characteristics of cognitive samples selected by our RAP method and (b) Comparison of the reasoning processes between our RAP method and the state-of-the-art LIMR [9].
@@
IV-C Key Insights and Discussion
Effect of RAP on reasoning upper bound. As illustrated in Fig. 7(a), our RAP method converges faster than the baseline, achieving optimal performance within 40 training steps compared to 100 for the baseline. These results demonstrate RAP’s efficiency in enhancing multi-modal reasoning while reducing training overhead and data redundancy. However, further analysis reveals differences between the full RAP model and its variant using only ACE and CDE, particularly during later training stages. This occurs because ACE and CDE inevitably retain easy examples, while discarding challenging yet informative samples. For example, multi-modal examples predicted correctly in only one out of five attempts but consistently mispredicted under text-only conditions may be mistakenly eliminated when the discrepancy threshold $\lambda_{c}$ exceeds 0.2. Such filtering reduces dataset complexity, limiting the model’s overall reasoning upper bound. To address this, we propose the Difficulty-aware Replacement Module (DRM) to explicitly replace easy samples with these informative yet challenging examples, thus elevating reasoning performance in later training stages.